OpenStudy (anonymous):

An airplane over the Pacific Ocean sights an atoll at an 7° angle of depression. If the plane is 663 m above the water, how many kilometers is it from a point 663 m directly above the atoll?

OpenStudy (wolf1728):

tan (7°) = 663 / distance distance = 663 / tan (7°) distance = 663 / 0.12278 distance = 5,399.9022642124 So atoll distance is about 5.4 kilometers distant